# Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study

### v1

@@ III-A RQ1 (Vulnerability detection)
For C++, both GPT-4 Turbo and Gemini 1.5 Pro showed a solid performance, both achieving the highest $F_{1}$ score of 0.88 using VD-Sys1 + VD-User1. Particularly, GPT-4o surpassed other models in recall (1.0) across all prompt sets, which is similar to C.
In JavaScript, Gemini 1.5 Pro maintained a balance between precision (0.87) and recall (0.90) in VD-Sys1 + VD-User1; thus, it achieved the highest $F_{1}$ score (0.88). Although GPT-4o obtained the highest recall of 1.0 using two prompt sets, CodeLlama-7b, a model more than 10 times smaller in size, also reached a recall of 1.0 (VD-Sys2 + VD-User1).
Lastly, regarding the result in Java, GPT-4 Turbo and GPT-4o showed the best $F_{1}$ score of 0.85 using VD-Sys1 + VD-User1 used by GPt-4o and VD-Sys2 + VD-User2 used by GPT-4 Turbo. While other models did not achieved satisfactory performance, GPT-4 Turbo exhibited the highest precision across all prompt sets. In particular, VD-Sys1 + VD-User1, the highest precision of 0.86 (GPT-4 Turbo), was approximately 15% better than the second highest one (GPT-4o).
RQ1 (vulnerability detection): Overall, GPT-4 Turbo and GPT-4o achieved the best vulnerability detection scores across multiple languages ($F_{1}=0.81$ and $F_{1}=0.83$). GPT-4 Turbo was the most effective for minimizing false alarms. GPT-4 Turbo achieved the highest $F_{1}$ score in Python. GPT-4o demonstrated exceptional recall, particularly in C and C++. Gemini 1.5 Pro performed the best in JavaScript. GPT-4-based models were the most effective LLMs in Java.
III-B RQ2 (CWE classification)
Table V (CWE classification) presents the effectiveness results across the different prompt setups. Unlike the detection task, the scores vary significantly across configurations. Similarly to the detection task, overall, GPT-4 Turbo and GPT-4o showed the best multi-class classification scores. Classification demonstrates a more difficult task than classification, as GPT-4 Turbo achieved an $F_{1}$ score of $0.53$ and GPT-4o achieved an $F_{1}$ score of $0.63$. The best results are obtained using the few-shot learning technique and a generic role (CWE-Sys2 + CWE-UserF configuration).
@@ III-B RQ2 (CWE classification)
For Python, Gemini 1.5 Pro was the most effective model using CWE-Sys2 + CWE-UserF setting. This exhibited the highest accuracy ($>$ 0.70) and $F_{1}$ (0.62).
@@
RQ2 (CWE classification): GPT-4o achieved the best vulnerability classification scores across multiple languages ($recall=0.74$ and $F_{1}=0.63$). GPT-4 Turbo showed strong performance in C and JavaScript. Gemini 1.5 Pro outperformed in Python. Few-shot prompts consistently outperformed zero-shot approaches across all models and languages.
IV CodeGuardian
To help engineers understand potential vulnerabilities in source code within their development workflow, we developed CodeGuardian, a VSCode extension that features just-in-time LLM-powered vulnerability scanning. It supports developers by providing real-time vulnerability analysis of their source code. More specifically, a user can highlight code snippets in editor, and CodeGuardian provides instant vulnerability analysis of the highlighted code. This feature helps developers identify and address security flaws at a snippet level. Multi-snippet or multi-file analyses are not supported.
Our study identified GPT-4 Turbo and GPT-4o as the best models for CWE detection and classification (see Section III). Consequently, CodeGuardian provides users with the option to select these models, using Azure OpenAI Studio APIs in the back end. The back-end system receives the highlighted code from a user and processes it to be concatenated to a specific prompt. After the vulnerability analysis, the results are shown on the extension panel.
@@ V-A Experiment Design
Participants were divided into two groups based on their responses: an experimental group (using CodeGuardian) and a control group (using only Internet search). Each group consists of 11 participants. We mainly referred to each participant’s years of experience in software development and cybersecurity to allocate them into equally distributed groups.
Challenges Design. All participants were required to perform a set of challenges during the experiment, which was identifying possible software vulnerabilities. In this experiment, a total of six challenges across five programming languages were presented to both groups. Table VI shows all the code snippets used for this user study, each snippet contains a specific vulnerability. While five snippets were vulnerable, one of the Python codes was used as a Placebo (i.e., not vulnerable).
Table VI: An overview of the six challenges of our user study.
@@ 6 Python CWE-79: Cross-site Scripting
Independent Variable. The independent variable was the type of tool used by the participants: CodeGuardian (LLM-based) and Internet search/own knowledge.
@@
Experimental Procedure. The first author of this study demonstrated how to use CodeGuardian exclusively for the experimental group, taking approximately 2-3 minutes to explain the usage instructions. Each participant then evaluated six code snippets, identifying any vulnerabilities and naming them or providing their CWE labels. The experimental group used CodeGuardian to address these challenges, while the control group could only use Internet resources, except for any LLM-based application. Participants had a maximum of 5 minutes to answer each question; if they finished early, the session progressed to the next task. All participants worked on a dedicated workstation to ensure consistent environmental conditions for all participants. This workstation was an EC2 instance running Debian 12 OS, accessible to all participants via a browser-based VNC.
Qualitative Analysis. The participants in the experimental group were asked to answer a post-questionnaire for the qualitative analysis of CodeGuardian. The questions focused on assessing the user experience and effectiveness of our plugin.
Data Analysis. To compare dependent variables across the experimental group and the control group, we analyzed the results using the Mann-Whitney U test [39] (with $\alpha=0.05$) and the magnitude of the differences using the rank-biserial correlation to calculate the effect size [40]. We selected this non-parametric test as completion time and accuracy data were not normally distributed.
V-B RQ3 (usefulness)
Figure 1: Challenge completion accuracy between the experimental group (green) and control group (yellow).
V-C Completion Accuracy
Accuracy was calculated based on the amount of correct answers out of all six challenges. Figure 1 visualizes the boxplot of challenge completion accuracy for both groups. We used a one-sided research hypothesis since we assumed that the experimental group had a higher percentage of correct answers than the control group. Therefore, the null and alternative hypotheses were defined as follows:
  • $H_{0}$: The two groups are the same in terms of the total challenge completion accuracy.
  • $H_{1}$: Total challenge completion accuracy in the experiment group was higher than in the control group.
@@ V-C Completion Accuracy
V-D Completion Time
We measured the total challenge completion time (minutes:seconds) and compared the results statistically between the experimental and control groups. A one-sided research hypothesis was used for this test, as we hypothesized that the experimental group would complete all the challenges faster than the control group. Hence, the null and alternative hypotheses were defined as follows:
  • $H_{0}$: The two groups are the same in terms of the total challenge completion time.
  • $H_{1}$: Total challenge completion time in the experiment group is less than in the control group.
Figure 2 depicts the boxplots of completion time over all the challenges. The median and standard deviation in the experimental group were approximately 10:03 (mm:ss), and the standard deviation was 4:29 (mm:ss). On the contrary, the control group exhibited a mean of 24:59 (mm:ss) and a standard deviation of 4:58 (mm:ss). Although the standard deviation of both groups was almost identical, the experimental group completed all the challenges in 66% less time than the control group on average (around 15 minutes, on average).
The results of the Mann-Whitney U test showed a statistically significant difference (p-value = 0.00018) with a large effect size (0.95). Therefore, we reject the null hypothesis and conclude that the experimental group completed the challenges faster than the control group.
@@ V-D Completion Time
Figure 2: Challenge completion time between the experimental group (green) and control group (yellow).
Furthermore, Figure 3 illustrates the completion time per challenge. The median completion time in the experimental group was less than that of the control group in all challenges. Moreover, challenge 5 had the largest difference in performance (completion time) between the experimental and control groups by approximately 3.5 minutes. In contrast, challenge 2 in the experimental group showed a slight median difference compared to the control group($M_{exp}$=2:14 and $M_{ctl}$=3:20). Additionally, the lowest observed standard deviation for the experimental group was 46 seconds (Challenge 4), whereas the highest was 1 minute and 13 seconds (Challenge 1). Similar to the experimental group, the lowest observed median and standard deviation for the control group were 3:20 (Challenge 2) and 50 seconds (Challenge 6), and the highest median and standard deviation were 5 minutes (Challenge 1 and 6) and 1 minute and 34 seconds (Challenge 1). Although there were a few outliers in both experimental and control groups, the gaps between the lowest and highest standard deviations were relatively similar. Additionally, the experimental group completed the challenges on average by approximately 47% (Challenge 1), 44% (challenge 2), 58% (challenge 3), 59% (challenge 4), 65% (challenge 5), and 55% (challenge 6) faster, compared to the control group.
Figure 3: Comparison of challenge completion time per challenge between the experimental group (green) and control group (yellow).
RQ3 (usefulness): Both statistical tests for the challenge completion time and accuracy showed statistically significant improvements in the experimental group compared to the control group. CodeGuardian successfully increased the speed and accuracy of completing the challenges with large effect size.
VI Threats to Validity
VI-1 Internal validity
We compared all LLMs under identical experimental settings and on the same benchmark. Our implementation of the testing scripts to evaluate the scores was tested thoroughly. Moreover, we evaluated the LLMs with a limited set of prompts.
Our user study involved 22 participants, a reasonable number due to selection constraints (i.e., prior industrial knowledge, experience in software development, and cybersecurity). Although the sample size was small, it achieved the necessary statistical test power. The varying levels of software development and cybersecurity expertise among our participants, including their familiarity with programming languages and VSCode, could have influenced our results.
To mitigate this threat, we collected relevant background information through a pre-questionnaire and divided the participants into two groups, striving for an equal distribution of expertise. Also, we used the same environment for all participants (a workstation with Debian 12 OS and a connection speed of around 700 Mbps). However, the response time of the GPT-4 Turbo API was occasionally slower, especially during periods of high usage due to the limited number of regions where the model is hosted on Azure OpenAI Studio. The slower response times could potentially impact user experience and the seamless integration of CodeGuardian into the development workflow. These issues were reflected in the users’ responses to the post-questionnaire. However, it is important to note that this is a limitation of the current hosting infrastructure on Azure OpenAI Studio.
VI-2 External validity
The limited number of LLMs in our evaluation poses a threat in terms of the generalizability of our results, especially with the availability of newer models. Our results may not generalize or generalize differently when considering other LLMs. To mitigate this issue, we selected the best models available as of 05.2024. Likewise, our paper only involved five programming languages (C, C++, Python, Java, and JavaScript). This limited scope might not cover the diversity of programming environments and their associated vulnerabilities.
Furthermore, we used a relatively small number of code challenges in our user study. This was due to the limited availability of our participants, most of whom were industry professionals with constrained time. As a result, we included only one snippet per language, with an additional placebo snippet for Python. While this approach provided valuable insights into the effectiveness of CodeGuardian for vulnerability analysis, it may not fully represent all vulnerabilities and coding scenarios encountered in real-world projects.
VII Discussion
VII-A Model Selection
GPT-4 Turbo often exhibited higher effectiveness compared to the other large language models (see Section III). Although GPT-4o is the successor to GPT-4 Turbo, it did not consistently outperform its predecessor. GPT-4o achieved higher recall across all prompt setups, indicating it could identify more true positives. However, overall, GPT-4 Turbo excelled in three specific prompt setups (VD-Sys1 + VD-User2, VD-Sys2 + VD-User1, and VD-Sys2 + VD-User) primarily due to GPT-4o’s lower precision. This highlights a crucial aspect of using LLMs for vulnerability detection: newer versions or successors do not always outperform older models. Therefore, model selection should be based on specific security priorities. For instance, GPT-4o is preferable for detecting all potential vulnerabilities (i.e., reducing false negatives), whereas GPT-4 Turbo is better for minimizing false alarms. This underscores the importance of selecting the right model based on the particular requirements.
VII-B Model Performance in Each Language
@@
Our study revealed significant variations in LLM performance across different programming languages for both vulnerability detection and CWE classification tasks. GPT-4-based models mostly outperformed other models across all languages. However, the strong performance of Gemini 1.5 Pro in JavaScript (vulnerability detection) and Python (CWE classification) indicates that different models may have strengths in particular languages or types of code analysis. This underscores the importance of considering language-specific performance when we select for secure code review.
VII-C Complexity of CWE Classification using LLM
Compared to vulnerability detection, our experiments reveal significant challenges in CWE classification, with performance varying across different CWE classes and model sizes. Larger models like GPT-4o, GPT-4 Turbo, and Gemini 1.5 Pro consistently outperformed smaller ones, particularly when employing few-shot learning techniques. For instance, we observed that recall rates of GPT-4o approximately triple using the few-shot setting. This improvement suggests that providing relevant examples helps models better understand the classification task. However, using few-shot learning does not help for smaller models, such as CodeLlama models due to the limited parameter size restricting their ability to effectively utilize the few-shot examples.
VII-D User Study
In our user study, the experimental group consistently completed challenges faster than the control group. Specifically, our placebo code analysis (Challenge 5) showed that most participants who used CodeGuardian were able to answer in under approximately two minutes. However, the control group’s participants took, on average, almost all of the allotted time (5 minutes). Thus, CodeGuardian helped participants quickly identify whether the code was vulnerable or not. We believe that this effect is because of the two-stage processing of CodeGuardian (see Section IV).
Furthermore, in Challenge 4, the most considerable reduction in completion time was found when participants used CodeGuardian. Analysis of our pre-questionnaire revealed that the participants were less skilled in JavaScript than in popular programming languages, such as Java or Python. Hence, we assume that, although some participants are not familiar with certain languages, CodeGuardian enables developers to analyze code and identify potential vulnerabilities regardless of prior knowledge.
In addition, accuracy was notably higher in the experimental group than in the control group. Evidence suggests that seamless, instant feedback and automated vulnerability analysis by CodeGuardian not only accelerate the analysis process but also correct vulnerability identification. Despite the majority of participants having limited knowledge of cybersecurity, according to the pre-questionnaire, the experimental group had an accuracy of more than 0.90 on average. In contrast, the accuracy of the control group was less than 0.50. This indicates that CodeGuardian effectively supports developers in identifying potential software vulnerabilities without having prior cybersecurity knowledge. However, our tool requires strict usage guidelines for junior developers, who might overly depend on their outputs without proper verification. Indeed, LLMs are prone to hallucinations, where the output appears plausible but is largely incorrect.
We believe CodeGuardian could have significant implications for secure software development practices. Reducing the time and effort required to identify vulnerabilities allows developers to conduct security checks more seamlessly within their workflow, potentially leading to more secure code being produced at a faster rate. The feedback from the experimental group indicated that the majority of participants ($>$ 70%) were satisfied with the speed of the vulnerability scanning process. Furthermore, users particularly appreciated how CodeGuardian effectively checked potential vulnerabilities automatically.
VIII Related Work
VIII-A Conventional Methods and Tools
Conventional vulnerability analysis primarily relies on static application security testing (SAST) and dynamic application security testing (DAST). SAST examines source code without execution and focuses on syntactic and semantic checks. DAST evaluates applications during runtime, mimicking attacker behavior [41]. However, SAST typically generates a high number of false positives [42], and DAST provides high false negatives, which may miss many vulnerabilities [43].
In constrast, in this paper we focus on LLM-based approaches to vulnerability detection.
VIII-B ML and DL in Vulnerability Detection
Recently, vulnerability detection systems have become more adapted to learning from data to predict and identify vulnerabilities with higher accuracy [16]. DL approaches, particularly utilizing Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, have further improved detection accuracy without requiring feature engineering by hand [15, 44]. Despite the advancement brought by DL, these models treat code as linear sequences, potentially overseeing complex semantic relationships as well as long-range dependencies [45].
These approaches are challenged by the development of pre-trained LLMs. Unlike DL approaches, LLMs such as the ones used in this paper require no explicit training, while achieving high scores in vulnerability detection.
VIII-C LLMs in Vulnerability Detection and Classification
Studies have shown that fine-tuned models like BERT and GPT outperform traditional DL models in vulnerability detection [10, 9]. However, this approach requires substantial computational resources and large, well-annotated datasets. In contrast, prompt engineering leverages pre-trained LLMs without any parameter update, offering a more resource-efficient approach. Several studies have demonstrated the effectiveness of prompt engineering, such as few-shot learning and chain-of-thought in vulnerability detection and classification [46, 47, 48]. Specifically, Zhou et al. [37] showed that GPT-4, using prompt engineering, outperformed fine-tuned CodeBERT.
Although prompt engineering does not require a large dataset, the effectiveness of LLM in vulnerability detection and classification heavily relies on the quality of the dataset used for evaluation. Recent studies highlighted data quality issues in existing datasets, such as poor data quality, low label accuracy, and high data duplication rates [26, 27], causing misleading results when measuring the effectiveness of LLMs.
Since this study focused on optimizing input prompts and inference methods without fine-tuning, it was important to ensure the use of a high-quality dataset. Moreover, to the best of our knowledge, existing works are limited in terms of language diversity, which is a characterizing element of our work.
VIII-D Tools Leveraging LLMs for Vulnerability Analysis
Although ChatGPT is not specifically designed to analyze and detect vulnerabilities, it has shown promising results [49, 14]. However, it is challenging to integrate ChatGPT seamlessly into the development workflow. To address this limitation, researchers have also developed IDE-integrated solutions [50, 51]. Yet, current tools rely on less advanced LLMs, potentially limiting their effectiveness compared to state-of-the-art LLMs like GPT-4.
To this aim, we developed CodeGuardian as a widely accessible VSCode extension that makes state of the art LLM models accessible. Our user study with industrial participants suggested the benefits and perils of a more pervasive adoption of LLMs for vulnerability detection beyond purely academic studies, a novel contribution to the state of the art.
IX Conclusions and Future Work
This paper presented a comprehensive study on the effectiveness of large language models (LLMs) for secure code assessment across multiple programming languages. We evaluated six state-of-the-art LLMs on vulnerability detection and CWE classification tasks using our developed multi-language vulnerability dataset. Our results demonstrate that GPT-4 Turbo and GPT-4o outperform other LLMs. Although the GPT-4-based models often achieved the highest effectiveness in both tasks and different programming languages, Gemini 1.5 Pro showed strong performance in JavaScript and Python.
Furthermore, we found that few-shot learning techniques significantly improve performance, especially for CWE classification. To bridge the gap between research findings and practical applications, we developed CodeGuardian integrating LLM-powered vulnerability analysis into developers’ workflows. Our user study with 22 industry-experienced developers showed that CodeGuardian significantly improves both the speed and accuracy of vulnerability detection, with participants completing tasks 66% faster and achieving 203% higher accuracy compared to traditional methods.
For future work, we should focus on expanding our manually annotated dataset to improve robustness and generalizability. Exploring advanced prompt engineering methods such as chain-of-thought and Retrieval Augmented Generation (RAG) could further enhance LLM performance. Additionally, improving CodeGuardian’s user interface and incorporating user feedback will make it a more effective tool for secure software development.

### v2

@@ 3.1. RQ1 (Vulnerability detection)
For C++, both GPT-4 Turbo and Gemini 1.5 Pro showed a solid performance, both achieving the highest $F_{1}$ score of 0.88 using Sys1–U1. Particularly, GPT-4o surpassed other models in recall (1.0) across all prompt sets, which is similar to C.
In JavaScript, Gemini 1.5 Pro maintained a balance between precision (0.87) and recall (0.90) in Sys1–U1; thus, it achieved the highest $F_{1}$ score (0.88). Although GPT-4o obtained the highest recall of 1.0 using two prompt sets, CodeLlama-7b, a model more than 10 times smaller in size, also reached a recall of 1.0 (VD-Sys2 + U1).
Lastly, regarding the result in Java, GPT-4 Turbo and GPT-4o showed the best $F_{1}$ score of 0.85 using Sys1–U1 used by GPt-4o and VD-Sys2 + VD-User2 used by GPT-4 Turbo. While other models did not achieved satisfactory performance, GPT-4 Turbo exhibited the highest precision across all prompt sets. In particular, Sys1–U1, the highest precision of 0.86 (GPT-4 Turbo), was approximately 15% better than the second highest one (GPT-4o).
RQ1 (vulnerability detection): Overall, GPT-4 Turbo and GPT-4o achieved the best vulnerability detection scores across multiple languages ($F_{1}=0.81$ and $F_{1}=0.83$). GPT-4 Turbo was the most effective for minimizing false alarms. GPT-4 Turbo achieved the highest $F_{1}$ score in Python. GPT-4o demonstrated exceptional recall, particularly in C and C++. Gemini 1.5 Pro performed the best in JavaScript. GPT-4-based models were most effective in Java.
3.2. RQ2 (CWE classification)
Table 6 (CWE classification) presents the effectiveness results across the different prompt setups. Unlike the detection task, the scores vary significantly across configurations. Similarly to the detection task, overall, GPT-4 Turbo and GPT-4o showed the best multi-class classification scores. Classification demonstrates a more difficult task than classification, as GPT-4 Turbo achieved an $F_{1}$ score of $0.53$ and GPT-4o achieved an $F_{1}$ score of $0.63$. The best results are obtained using the few-shot learning technique and a generic role (CWE-Sys2 + CWE-UserF configuration).
@@ 3.2. RQ2 (CWE classification)
For Python, Gemini 1.5 Pro was the most effective model using CWE-Sys2 + CWE-UserF setting. This exhibited the highest accuracy (>> 0.70) and $F_{1}$ (0.62).
@@
RQ2 (CWE classification): GPT-4o achieved the best overall vulnerability classification performance across languages ($recall=0.74$, $F_{1}=0.63$). GPT-4 Turbo performed best in C and JavaScript, while Gemini 1.5 Pro outperformed in Python. Across all models and languages, few-shot prompting consistently outperformed zero-shot.
4. User Study
4.1. CodeGuardian
To help engineers understand potential vulnerabilities in source code within their development workflow, we developed CodeGuardian, a VSCode extension that features just-in-time LLM-powered vulnerability scanning. It supports developers by providing real-time vulnerability analysis of their source code. More specifically, a user can highlight code snippets in editor, and CodeGuardian provides instant vulnerability analysis of the highlighted code. This feature helps developers identify security flaws at a snippet level. Multi-snippet or multi-file analyses are not supported.
Our study identified GPT-4 Turbo and GPT-4o as the best models for CWE detection and classification (see Section 3). Consequently, CodeGuardian provides users with the option to select these models, using Azure OpenAI Studio APIs in the back end. The back-end system receives the highlighted code from a user and processes it to be concatenated to a specific prompt. After the vulnerability analysis, the results are shown on the extension panel. Additional details of CodeGuardian are available in our replication package (rep-package, 2025).
@@ 4.2. Experiment Design
Table 7. An overview of the six challenges of our user study.
@@ 6 Python CWE-79: Cross-site Scripting
Participants were divided into two groups based on their responses: an experimental group (using CodeGuardian) and a control group (using only Internet search). Each group consists of 11 participants. By balancing expertise in software development and cybersecurity, we ensured that the experimental and control groups had comparable backgrounds.
We mainly referred to each participant’s years of experience in software development and cybersecurity to allocate them into equally distributed groups. Lastly, none of the participants had any experience with CodeGuardian.
Challenges Design. All participants were required to perform a set of challenges during the experiment, which was identifying possible software vulnerabilities. In this experiment, a total of six challenges across five programming languages were presented to both groups. Table 7 shows all the code snippets used for this user study, each snippet contains a specific vulnerability. While five snippets were vulnerable, one of the Python codes was used as a Placebo (i.e., not vulnerable).
Independent Variable. The independent variable was the type of tool used by the participants: CodeGuardian (LLM-based) and Internet search/own knowledge. More specifically, Level 1 expects participants to use CodeGuardian to identify vulnerabilities, and Level 2 expects them to use their knowledge.
@@
Experimental Procedure. The first author of this study demonstrated how to use CodeGuardian exclusively for the experimental group, taking approximately 2-3 minutes to explain the usage instructions. Each participant then evaluated six code snippets, identifying any vulnerabilities and naming them or providing their CWE labels. The experimental group used CodeGuardian to address these challenges, while the control group could only use Internet resources, except for any LLM-based application. Participants had a maximum of 5 minutes to answer each question; if they finished early, the session progressed to the next task.
All participants worked on a dedicated workstation to ensure consistent environmental conditions for all participants. This workstation was an EC2 instance running Debian 12 OS, accessible to all participants via a browser-based VNC.
Data Analysis. To compare dependent variables across the experimental group and the control group, we analyzed the results using the Mann-Whitney U test (Mann and Whitney, 1947) (with $\alpha=0.05$) and the magnitude of the differences using the rank-biserial correlation to calculate the effect size. We selected this non-parametric test as the completion time and accuracy data were not normally distributed.
4.3. RQ3 (usefulness)
Figure 1. Challenge completion accuracy between the experimental group (green) and control group (yellow).
4.3.1. Completion Accuracy
Accuracy was calculated based on the amount of correct answers out of all six challenges. Figure 1 visualizes the boxplot of challenge completion accuracy for both groups. We used a one-sided research hypothesis since we assumed that the experimental group had a higher percentage of correct answers than the control group. Therefore, the null and alternative hypotheses were defined as follows: $H_{0}$: The two groups are the same in terms of the total challenge completion accuracy; $H_{1}$: Total challenge completion accuracy in the experiment group was higher than in the control group.
@@ 4.3.1. Completion Accuracy
Figure 2. Challenge completion time between the experimental group (green) and control group (yellow).
4.3.2. Completion Time
We measured the total challenge completion time (minutes:seconds) and compared the results statistically between the experimental and control groups. A one-sided research hypothesis was used for this test, as we hypothesized that the experimental group would complete all the challenges faster than the control group. Hence, the null and alternative hypotheses were defined as follows: $H_{0}$: The two groups are the same in terms of the total challenge completion time; $H_{1}$: Total challenge completion time in the experiment group is less than in the control group.
Figure 2 depicts the boxplots of completion time over all the challenges. The median and standard deviation in the experimental group were approximately 10:03 (mm:ss), and the standard deviation was 4:29 (mm:ss). On the contrary, the control group exhibited a mean of 24:59 (mm:ss) and a standard deviation of 4:58 (mm:ss). Although the standard deviation of both groups was almost identical, the experimental group completed all the challenges in 66% less time than the control group on average (around 15 minutes, on average). The results of the Mann-Whitney U test showed a statistically significant difference (p-value = 0.00018) with a large effect size (0.95). Therefore, we reject the null hypothesis and conclude that the experimental group completed the challenges faster than the control group.
RQ3 (usefulness): CodeGuardian significantly improved speed and accuracy, with large effect sizes.
5. Threats to Validity
Internal validity. All LLMs were compared under identical settings and on the same benchmark. Our testing scripts were validated, but the limited prompt set may affect generalizability. We believe the selected models are representative. The user study involved 22 participants with prior industrial experience in software development and cybersecurity. Although the sample size was small, it provided sufficient statistical power. Differences in participants’ expertise (e.g., programming languages, VSCode) may have influenced results. To mitigate this, we collected background data, balanced expertise across two groups, and ensured identical environments (Debian 12 workstation, 700 Mbps connection). Response times of GPT-4 Turbo occasionally varied due to Azure hosting limitations, which was reflected in post-questionnaire feedback.
External validity. The evaluation covered a limited set of LLMs and programming languages (C, C++, Python, Java, JavaScript). Results may differ for other models or languages. To reduce this threat, we used state-of-the-art models as of 05.2024. The user study also included a small number of code challenges (one per language plus a Python placebo), constrained by participants’ availability. While useful for evaluating CodeGuardian, this may not represent the full spectrum of vulnerabilities and real-world scenarios (Dozono et al., 2025).
6. Discussion
Model Selection. Although GPT-4 Turbo is the predecessor of GPT-4o, it often outperforms other models. While GPT-4o achieved higher recall across all prompt setups, its lower precision caused it to underperform in some cases. This suggests that newer LLM versions do not always surpass earlier ones. Model selection should thus reflect security priorities: GPT-4o is preferable when minimizing false negatives, whereas GPT-4 Turbo better reduces false positives.
@@ 6. Discussion
Performance Across Languages. LLM performance varied significantly by programming language. GPT-4 models generally achieved the best results, but Gemini 1.5 Pro showed strong performance in JavaScript (vulnerability detection) and Python (CWE classification). This highlights the importance of considering language-specific performance when using LLMs for secure code review.
Vulnerability Type Analysis. LLMs were especially effective in detecting injection-related vulnerabilities. Detection rates for CWE-89 (SQL Injection) and CWE-79 (Cross-site Scripting) exceeded 81% across languages. This strength is particularly relevant for web application security. The user study further confirmed this: in Challenge 6 (CWE-79), participants using CodeGuardian completed the task 55% faster than the control group.
CWE Classification Complexity. CWE classification proved more difficult than vulnerability detection. Larger models (GPT-4o, GPT-4 Turbo, Gemini 1.5 Pro) clearly outperformed smaller ones, especially with few-shot prompts, which tripled GPT-4o’s recall. In contrast, few-shot learning did not improve CodeLlama models, suggesting that smaller models cannot effectively leverage examples due to limited capacity.
User Study. The experimental group consistently completed tasks faster and with higher accuracy. For instance, in the placebo task (Challenge 5), most participants using CodeGuardian answered in under two minutes, compared to nearly five minutes for the control group. Accuracy gains were equally notable: the experimental group achieved over 0.90 on average, versus less than 0.50 in the control group. These results show that CodeGuardian accelerates analysis and improves detection, even for developers with limited security expertise. However, strict usage guidelines are needed for junior developers, as LLM outputs remain prone to hallucinations.
7. Conclusions and Future Work
We presented a comprehensive study on the effectiveness of LLMs for secure code assessment across multiple programming languages. We evaluated six state-of-the-art LLMs on vulnerability detection and CWE classification using a newly developed multi-language dataset. Our results show that GPT-4 Turbo and GPT-4o generally achieved the best overall performance, while Gemini 1.5 Pro performed strongly in JavaScript and Python. To bridge research and practice, we introduced CodeGuardian, a VSCode extension that integrates LLM-powered vulnerability analysis into developers’ workflows. In a user study with 22 industry professionals, CodeGuardian improved task completion speed by 66% and accuracy by 203% compared to traditional tools. Future work will focus on expanding and refining the annotated dataset to enhance robustness and generalizability, exploring advanced prompting techniques such as chain-of-thought and retrieval-augmented generation (RAG), other foundation models (Gao et al., 2025; Baresi et al., 2025; Maryam et al., 2025; Merabishvili et al., 2026; Weißl et al., 2025) and tasks such as natural language conversations (Giebisch et al., 2025; Sorokin et al., 2026; Habicht et al., 2025), and improving CodeGuardian usability through interface enhancements and developer feedback.